A motorboat takes two thirds as much time to travel ten miles downstream as it does to travel the same distance upstream. If the rate of the current is 5 mph, find the speed of the motorboat in still water.
Answer by jorel1380(3719)

10/x+5=2/3(10/x-5)
10/(x+5)=20/3(x-5)
10(3)(x-5)=20(x+5)
30x-150=20x+100
10x=250
x=25
The motorboat's speed in still water is 25 mph.